OpenStudy (anonymous):

If the principal P = $900, the interest I = $702, and time t = 6 years, find the following. (a) What is the rate? 45.5 . % (b) What is the future value?

OpenStudy (anonymous):

the rate is not 45.5 thats a typo

OpenStudy (anonymous):

i=prt but my answers keep coming up wrong

OpenStudy (anonymous):

is it compounding or simple interest

OpenStudy (anonymous):

ok so it is simple interest

OpenStudy (anonymous):

702=900*r*6

OpenStudy (anonymous):

\[\large{Interest = Principal\times{rate}\times{time}}\] let us assume rate of interest is r \[\large{$702 = $900\times{r}\times{6}}\] now solve for r

OpenStudy (anonymous):

702=5400r

OpenStudy (anonymous):

702/5400

OpenStudy (anonymous):

0.13

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

what about the future value

OpenStudy (anonymous):

Future value means total amount you will be having after 6 years

OpenStudy (anonymous):

a=p+i

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0.13=900+702

OpenStudy (anonymous):

no just \[\large{{Amount}= Principal + Interest Earned}\]

OpenStudy (anonymous):

oic lol been studying to long 900 + 702

OpenStudy (anonymous):

thanks so much 1702

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

add properly